Friday, November 29, 2002

Mostly off today. Put together the new fish tank. We think there is room for another 10 gallon tank on the same cart. O, what fun.

Lazy day, watched the Muppet Christmas Movie. Cute.
No time to blog! Tuesday was pretty uneventful, although very busy at work. Wednesday, Kelly kinda sick, I had to pick him up after the nice lunch with Heather. We went to La Tolteca.

In at work this morning to do backups. This is going to get settled soon. See the other blog on this.

More later.

Monday, November 25, 2002

Boy, am I tired. Got to bed just after one and the alarm goes at 6:30. I snoozed it for half an hour. I was in the office by 7:45. I feel like heck. Problems galore at work. Can't really get away.

Sunday, November 24, 2002

No blog since Thursday, how can that be? That was the 21st, now it's the last hour of the 24th.

Whew, where to start?

Friday, the 22nd was one ugly day. 14 hours of fun, fun. Purging directories and records in PBS. On the phone constantly with PBS getting ready for the upgrade.

Work is just too much for this blog, so I made another blog as a team and invited Kelsie. I don't know if she will like it. INI_Tech

Saturday, was pretty useless. I didn't get dress until about 5pm. Watched Enders live, Trevor and Tom burn. Very sad. I am still way behind on Ender mail.

We went to the store and got stuff for Kelly to work on the banjo-mando and then went to pet store to get new fish tank.

Today, we played for Hospice Festival of Trees with Mike. Very nice.

Then, we went over Tom's and had a great time. We found out that Jill grows butterfly KOI! Wow! Has ponds in backyard. What fun.

Then to work, backing up, etc. This is going to be a very late night and I will need to come in early tomorrow! Yikes.

Thursday, November 21, 2002

Smile! Kelsie and ME!

Made a new web page out of the gizmo pix of our luncheon, today.
Cambridge migration went pretty well. PBS locked up the database about 10:51am and were done about 2:15pm. I tested some stuff and talked a few users throught the new steps to log on. Problems remain with the telemarketing people. They cannot comprehend my instructions.

Problems were only spool still had to be setup. User setup had to be looked at. The login names in AIX on the Dover side HAD to agree with user names within Camb PBS. Me and Sis had some differences from Dover to Cambridge with our logins and Sandi and Doretha had never had an AIX login on the Dover side. There may be others, but I don't think so.
Sinus medicine is not working well.

Still waiting to bring Cambridge down for migration.
No time to blog yesterday. Was up until 2 the night before. Wed (yesterday) I came in for a couple of hours then went home did laundry and took a nap. Then, came in later to work a few more hours.

PBS is configuring the new drives, everything going well.

Cambridge database gets moved today, more on that later.

Had rehearsal at Mike's house last night, looking toward Festival of Trees for hospice at Ag Museum on Sunday at one. It will be lovely. Just Mike, Kelly and me. We will play lots of slow airs and lovely harp with flutes, fiddles, concertina and whistle.
(wrote Tammy, Mike & Sheila to let them know)

Tuesday, November 19, 2002

Going HOME. All done!

Got to work today at 1pm. Things seem very calm.

The new dishwasher works pretty good. Things are starting to improve on the housekeeping front.

I wore the Sheela and an angel pin. Hoping to do whatever I can to improve my chances with the hardware upgrade.

Jeff, the IBM guy, called to make arrangements for tonight and check to see that I have the part. I told him I didn't trust him or IBM and we need this to work this time (I did qualify it to say that I trusted him to install the parts, though, terrific, huh). Probably shot myself in the foot, but I don't like how I have been treated.

More on this later.

Monday, November 18, 2002

What a day! Work, work, work. I called and cancelled manicure for tomorrow. Then I chipped a nail. Damn. Too busy with IBM. Will I LIVE to see next week??

I have to figure out what to bring to employee luncheon on Thurs. I just don't know how my day will be. Can't be something from my kitchen. Perhaps some day!

Kel is home working on the dishwasher thing now. That will be so great!

Time for more sinus medicine. I should go to bed early.
Not feeling too good. I think I am getting some sinus-thing. Goes hand in glove with any upgrade, does it not??
Again, no time to blog, here's how it went Friday night, or 11/15 con't....

Friday Nov 15
9:00 back up complete.
IBM Jeff is here. He does not like the way the drives are stacked in the
big box (Norm). Opens little box for new part. Worries about cables and

9:20 with systom down, looks, thinks we might ALSO need adapter board.

9:25 calls IBM, we don't need board but their is a data transfer cable- goes
back plane to back plane (part #93h4340) Jeff talked through all the steps
of intallation with support to make sure that there would be no more
surprises. I hope that's true.

10:00 reassembles machine

10:15 we have login - comparinq hdisk to before, yes, all there, error log
good, only shows off, on, reboot.

10:25 Jeff leaves

10:30 fixed cron back for weekend.


Back to Today, Monday 11/18....

Conversation with Summit about the weekend and the parts. Joe tells me that IBM said the drives were in but waiting for cable. That's untrue, they are all sitting in box. He can work with IBM, they will come out tomorrow and make it right with parts that they are sending overnight today.

I asked him to send SSA adapter and connections, I want to have more than I need and send back what is not used this time, no excuses. I referred to comment made by Sean in August about perhaps needing more than one SSA adapter. I no longer trust what IBM says about this install. Joe is checking with his F50 expert there and will get back.

More on this later!

Friday, November 15, 2002

There was NO time yesterday to BLOG. Atex was here and Kelsie & I had to help them with research on conversion (likely) to Enterprise Retail. 11 hour day.

Started today at 9:00am, not at my best, early mornings are not my time of day.

I know it's mid-November but the heat is on in the building and it was 82 in my office yesterday, so today, I wore t-shirt and shorts with Berkie sandals!! It's 64 outside, look at the time, have I had lunch?? Another day from HECK.

Wednesday, November 13, 2002

The FF board tonight, need to scan photos and update web page, o my.

May have to have possum rehearsal at the office, may be on duty! That would be not so good but better than none! Late, sending Kel to bed.

I am trying to finish Two Towers, anyhow. 90nite.
O, I asked Jon about the CUTE thing from yesterday and he said it's something Mel sent him, we have to love it. It starts to make sense and I like it better now!
LL gave me a tenner toward the palm today!
Mike P. offered us tickets to Eileen Ivers at the Schwartz, we have a Possum gig! How funny!

Tuesday, November 12, 2002

Lorraine busted her palm pilot. Great!
Jon sends me a CUTE thing. What's up with that?? CUTE
The mac would have worked, but I negleceted to hit the publish button, well, duh.

I did not feel much like going to work today. Wanted to stay in bed with the spotty cat. O, the wind and rain. The last two days have knocked off all the pretty fall leaves, trees are starting to look naked. The tomato patch across the street is now level. Abandoned tomatoes are left there to rot, green and red. Heaving a big sigh, there.

Here we go round and round with the work stuff. Peg sent PBS a request for some help with IBM labor bill to come back out. Waiting to hear.

I bet I BLOG a lot today!

Monday, November 11, 2002

Trying the blog at home from the Mac.

Heard back from Summit, sent big emails to everyone. Going to cost us more money.

Kel called, we are going to lunch. Veggie BK, yum.

Still waiting to hear from Summit.

Hmmm, Dan (?) called from BlackOak! Ha! they missed the boat! They still want to send us a package. We have other servers, this might not hurt, I said okay.

Homefront: Still need to order the crock-o-gleam. Pampered Chef stuff came today. Need to take the kitchen back, fresh cooked stuff just in time for the holidays! Wow.
testing this!
working on the upgrade at work, logging the progress, I was reminded that I need to blog!! Here's that log...
-----Original Message-----
From: Jan
Sent: Sunday, November 10, 2002 9:30 PM
To: Techteam; Peggy Teal
Subject: CIRC Hardware BLOG 11/10/02

x Tedious detail ALERT! x
If you are sensitive, delete or file now!

Here's how it went today, Sunday, 11/10/02...

12:15 I arrive to do the daily backup manually.

12:35 pounded out cron, locking the databases, timing the backup...

2:10 the directories are backed up to the tape, 1:40 elapsed time

3:05 the verify is done and tape ejected, total for all steps about 2-1/2

3:40 Norm gets here. First, he installed all the new memory

4:30 problem with ALL the hard drives
Part #76h2697 for a 4g drive on the top row w/white face plate (in the #1

There are 4 on middle row v/ black face plate. On the middle row, there are
6 slots left to rgt (X= drive in slot, O= open)--
1 2 3 4 5 6

There is no boardfor drives on the bottom row.

At this time, we went with staying with all the old drives. NO NEW ONES IN
THE SERVER. 2 of the new drives face the right way. We can get Summitt to
send a back set (or 4?) that will fit overnight tomorrow. We can install,
turn key, open front and drives slide right in.

Need to hear from Summit and PBS about what parts are right and make a new
plan. I am hoping to get this poked in there and back on schedule no later
than the end of the day Wednesday.

5:10 memory problem. F50 will not come back up. Norm calls support. They
confirm that new memory board should have gone in original position and new
one (because it is full w/16)and the old one (it has 12 slots full only)
should qo into next position. Norm tries to switch. System cannot even
tell us which of the chips may be bad,reseatin chips.. Try again.

5:37 we have a login.

5:47 Running diag, Norm runs into trouble with F7 key. Could be emulation.
But we switch keybds w/ no effect. Calls IBM for help.

6:20 I asked Norm if we could telnet from Adv, since it is ibm3151
emulation, we did, F7 -- this worked (I don't know if this happened to us

6:25 We have two bad chips, Norm is taking these out and moving them all
back up in sequence.

8:25 At last the diag errors were gone, after Norm reseat the chips, moved
'em around and restarted the system about 8 times.

8:35 I have the system back, checked lpstat, unlocked db and checked
application and spool

8:38 calling PBS

8:55 Gary dialed in, will apply the multivolume to the databases that need
it, will be done before 10 and will unlock the databases for the users.
Gary verifies that we now have 3GB memory, yay!

9:20 Norm called back, wanted work order and contract #!

9:30 I am going home, remind me that the cron needs to be set back so it
will run tomorrow night.
Here's the first blog. Learning to make this easy, since it's free. And since it's just for me to read... I am trying to be sane about this!